OpenStudy (anonymous):

solve the IVP, y'=abs(y-1), y(0)=1

OpenStudy (raffle_snaffle):

is this diff eqs?

OpenStudy (anonymous):

yeah

OpenStudy (raffle_snaffle):

you mean calculus or differential equations?

OpenStudy (anonymous):

i know that abs(y-1)=sqrt((y-1)^2) but i cant see how to apply it and i have just started studying diff eq

OpenStudy (anonymous):

it is an initial value problem from a differential equations class

OpenStudy (raffle_snaffle):

have not taken the class yet, sorry

OpenStudy (anonymous):

Suppose that y>=1, then You can verify that \[ y= 1 - e^x \] is a solution

OpenStudy (anonymous):

Suppose that y<=1, then You can verify that \[ y= 1- e^{-x} \] is a solution.

OpenStudy (anonymous):

how does dy/dx = -e^x solve dy/dx=abs(y-1)??

OpenStudy (anonymous):

you mean e^x + 1 ?

OpenStudy (anonymous):

we suppose in each case that y >1 or y<1

OpenStudy (anonymous):

ok so walk me through this, sorry. if y = 1 - e^x and y >1 then x<ln(1) = 0 right? and then dy/dx = -e^x but - e^x is negative for this interval and so isnt a solution to the diff eq

OpenStudy (anonymous):

The first solution is not acceptable sinc \[ y = 1 - e^x <1 \] and the hypothesis was for y >1

OpenStudy (anonymous):

Let concentrate on y < 1, so y -1 <0 and Abs(y-1)=1-y

OpenStudy (anonymous):

i see how this solves the condition that y(0)=1 but do not see how it solves dy/dx=abs(y-1), it works if you ignore the abs but not otherwise

OpenStudy (anonymous):

ok i follow

OpenStudy (anonymous):

Use the initial condition to get c =0

OpenStudy (anonymous):

\[ \frac {dy}{dx}= 1- y\\ \frac {dy}{1-y}= dx\\ -\ln(1-y) = x +c\\ e^{-\ln(1-y)}= e^{x+c}\\ \frac 1 {1-y} =e^{x+c}\\ 1-y= e^{-x +c}\\ y = 1 - e^{-x -c}\\ \]

OpenStudy (anonymous):

very clear, thank you :)

OpenStudy (anonymous):

YW